OpenStudy (anonymous):

what value of c makes the polynomial below perfect square? 4x^2+12x+c

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

+4x%5E2 is the square of 2x

OpenStudy (anonymous):

make the co-efficient as 1 \[4x ^{2}+12x+c=4\left( x ^{2}+3x \right)+c\] \[ add \left( \frac{ co-efficient of x }{2 } \right)^{2}\] \[=4\left( x ^{2} +3x+\left( \frac{ 3 }{ 2 } \right)^{2}-\left( \frac{ 3 }{ 2 } \right)^{2 } \right)+c\] \[=4\left( x+\frac{ 3 }{ 2 } \right)^{2}-4*\frac{ 9 }{ 4 }+c\] \[=4\left( x+\frac{ 3 }{ 2} \right)^{2}+c-9\] If c-9=0 ,it is a perfect square. or c=9

OpenStudy (anonymous):

misprint, \[make the co-efficient of x ^{2} as 1.\]